OpenStudy (anonymous):

having a mental skid... how do you calculate by hand the derivative for Log [1/x] ?

OpenStudy (anonymous):

I know Log[x] --> 1/x but it cant be 1/(1x) can it?

jimthompson5910 (jim_thompson5910):

if the log is base e, ie a natural log, then log(1/x) = log(x^(-1)) = -1*log(x) then you would derive that to get -1/x

OpenStudy (anonymous):

yes thats the base .. ahh right on.. extract the -1 via the exponent.

OpenStudy (anonymous):

does it help to know that \[\log(\frac{1}{x})=-\log(x)\]

OpenStudy (anonymous):

yes they both help.. thanks... whats the name of that identity there satellite?

OpenStudy (dan815):

log power rule, log (x^a) = a*log(x)